# Inference-Time Alignment for Diffusion Models via Doob’s Matching

### v1

@@ Appendix B Derivations in Section 3
Appendix C Error Bounds for the Doob’s Guidance Estimation
C.1  The regularity of the Doob’s $h$-function
@@
C.2  Oracle inequality of the Doob’s $h$-function matching
@@
Suppose Assumptions 1 and 2 hold. Let $t\in(0,T)$ and let ${\mathscr{H}}_{t}$ be a convex hypothesis class. Let $\widehat{h}_{t}^{\lambda}$ be the gradient penalized empirical risk minimizer defined as (3.21), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequalities hold:
$\lesssim\underbrace{\inf_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{H^{1}(p_{T-t})}^{2}}_{\text{approximation}}+\underbrace{\bar{B}{\mathfrak{R}}_{n}({\mathscr{H}}_{t})+\frac{\lambda\bar{B}}{\sigma_{T-t}^{2}}{\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t})}_{\text{generalization}}+\underbrace{\frac{\lambda^{2}d^{2}\bar{B}^{2}}{\sigma_{T-t}^{8}}}_{\text{regularization}},$
$\lesssim\underbrace{\frac{1}{\lambda}\inf_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{H^{1}(p_{T-t})}^{2}}_{\text{approximation}}+\underbrace{\frac{\bar{B}}{\lambda}{\mathfrak{R}}_{n}({\mathscr{H}}_{t})+\frac{\bar{B}}{\sigma_{T-t}^{2}}{\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t})}_{\text{generalization}}+\underbrace{\frac{\lambda d^{2}\bar{B}^{2}}{\sigma_{T-t}^{8}}}_{\text{regularization}},$
@@
We first provide a proof sketch of this lemma.
  1. (i)
    In this step, we establish the following equality of the gradient penalized empirical risk minimizer $\widehat{h}_{t}^{\lambda}$ based on the variation inequality:
    (C.5) $\overbrace{{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}\Big]}^{\text{(I)}}+\lambda\overbrace{{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\|\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}\Big]}^{\text{(II)}}$
    $\leq\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+V_{t}^{2}$
    $\quad+2\lambda\underbrace{{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\langle\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i}),\nabla h_{t}(\mathbf{X}_{T-t}^{i})-\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})\rangle\Big]}_{\text{(III)}}.$
  2. (ii)
    In this step, we provide the generalization error bounds for terms (I) and (II) in (C.5):
    ${\mathbb{E}}\big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\big]+V_{t}^{2}$
    (C.6) $\leq{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}(\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-w(\mathbf{X}_{0}^{i}))^{2}\Big]+12\bar{B}{\mathfrak{R}}_{n}({\mathscr{H}}_{t}),$
    (C.7) $\leq{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\|\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})-\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i})\|_{2}^{2}\Big]+\frac{16\bar{B}}{\sigma_{T-t}^{2}}{\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t}).$
  3. (iii)
    In this step, we estimate term (III) in (C.5):
    (C.8) $\lambda{\mathbb{E}}\Big[\frac{1}{n}\sum_{i=1}^{n}\langle\nabla h_{t}^{*}(\mathbf{X}_{T-t}^{i}),\nabla h_{t}(\mathbf{X}_{T-t}^{i})-\nabla\widehat{h}_{t}^{\lambda}(\mathbf{X}_{T-t}^{i})\rangle\Big]-\frac{1}{2}{\mathbb{E}}\big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\big]$

### v2

@@ Appendix B Derivations in Section 4
Lemma B.1.
Suppose Assumptions 1 and 2 hold. Assume that $v_{t}\in H^{1}(p_{T-t})$. Then
$(\nabla h_{t}^{*},\nabla v_{t})_{L^{2}(p_{T-t})}=(\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t},v_{t})_{L^{2}(p_{T-t})}.$
Proof of Lemma B.1.
We first construct a sequence of cut-off functions $\{\psi_{k}\}_{k=1}^{\infty}\subseteq C_{0}^{\infty}({\mathbb{R}}^{d})$, satisfying
  1. (i)
    $\psi_{k}(\mathbf{x})=1$ for $\|\mathbf{x}\|_{2}\leq k$,
  2. (ii)
    $\psi_{k}(\mathbf{x})=0$ for $\|\mathbf{x}\|_{2}\geq 2k$,
  3. (iii)
    $\psi_{k}(\mathbf{x})\in(0,1)$ for $\mathbf{x}\in{\mathbb{R}}^{d}$, and
  4. (iv)
    $\|\nabla\psi_{k}(\mathbf{x})\|_{2}\leq Ck^{-1}$ for some constant $C$ independent of $\mathbf{x}$ and $k$.
See Brezis [2011, Theorem 8.7] for a detailed construction of such cut-off functions. Then we focus on the compactly supported approximations $\{\psi_{k}v_{t}\}_{k=1}^{\infty}$:
$-\langle\nabla h_{t}^{*},\nabla(\psi_{k}v_{t})\rangle_{L^{2}(p_{T-t})}$
$=-\int_{{\mathbb{R}}^{d}}\langle\nabla h_{t}^{*}(\mathbf{x}),\nabla(\psi_{k}v_{t})(\mathbf{x})\rangle p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(B.4) $=-\int_{{\mathbb{R}}^{d}}\nabla\cdot(\nabla h_{t}^{*}(\mathbf{x})p_{T-t}(\mathbf{x})(\psi_{k}v_{t})(\mathbf{x}))\,\mathrm{d}\mathbf{x}+\int_{{\mathbb{R}}^{d}}\nabla\cdot(\nabla h_{t}^{*}(\mathbf{x})p_{T-t}(\mathbf{x}))(\psi_{k}v_{t})(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(B.5) $=\int_{{\mathbb{R}}^{d}}\nabla\cdot(\nabla h_{t}^{*}(\mathbf{x})p_{T-t}(\mathbf{x}))(\psi_{k}v_{t})(\mathbf{x})\,\mathrm{d}\mathbf{x},$
where we used the Gauss’s divergence theorem and the fact that $\psi_{k}v_{t}\in H_{0}^{1}(B({\mathbf{0}},k))$ with $B({\mathbf{0}},k)\coloneq\{\mathbf{x}:\|\mathbf{x}\|_{2}\leq k\}$. For the left-hand side of (B.5), we have
$\int_{{\mathbb{R}}^{d}}\langle\nabla h_{t}^{*}(\mathbf{x}),\nabla(\psi_{k}v_{t})(\mathbf{x})\rangle p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(B.6) $=\int_{{\mathbb{R}}^{d}}\psi_{k}(\mathbf{x})\langle\nabla h_{t}^{*}(\mathbf{x}),\nabla v_{t}(\mathbf{x})\rangle p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}+\int_{{\mathbb{R}}^{d}}\langle\nabla h_{t}^{*}(\mathbf{x}),\nabla\psi_{k}(\mathbf{x})\rangle v_{t}(\mathbf{x})p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}.$
For the second term in (B.6), we have
$\Big|\int_{{\mathbb{R}}^{d}}\langle\nabla h_{t}^{*}(\mathbf{x}),\nabla\psi_{k}(\mathbf{x})\rangle v_{t}(\mathbf{x})p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}\Big|$
$\leq\int_{{\mathbb{R}}^{d}}\|\nabla h_{t}^{*}(\mathbf{x})\|_{2}\|\nabla\psi_{k}(\mathbf{x})\|_{2}|v_{t}(\mathbf{x})|p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
$\leq\frac{C}{k}\int_{{\mathbb{R}}^{d}}\|\nabla h_{t}^{*}(\mathbf{x})\|_{2}|v_{t}(\mathbf{x})|p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
$\leq\frac{C}{k}\|\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}\|v_{t}\|_{L^{2}(p_{T-t})},$
where the second inequality holds from the definition of the cut-off function, and the last inequality is due to Cauchy-Schwarz inequality and the fact that $h_{t}^{*}\in H^{1}(p_{T-t})$, which is a direct conclusion of Lemmas C.1 and C.4. Taking limitation with respect to $k\to\infty$ and using Lebesgue’s dominated convergence theorem yields
(B.7) $\lim_{k\to\infty}\int_{{\mathbb{R}}^{d}}\langle\nabla h_{t}^{*}(\mathbf{x}),\nabla\psi_{k}(\mathbf{x})\rangle v_{t}(\mathbf{x})p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}=0.$
Combining (B.5), (B.6), and (B.7) and taking limitation with respect to $k\to\infty$ completes the proof. ∎
Proposition 4.2.
Let $\lambda>0$, $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10), and $h_{t}^{\lambda}$ be the minimizer of ${\mathcal{J}}_{t}^{\lambda}$ defined as (4.3). Then $h_{t}^{*}\in H^{2}(p_{T-t})$, and
$\|h_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq\lambda^{2}\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2},$
$\|\nabla h_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq\lambda\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2}.$
Proof of Proposition 4.2.
First, $h_{t}^{*}\in H^{2}(p_{T-t})$ is a direct conclusion of Lemmas C.1, C.4, and C.5. It remains to prove two inequalities. Using Proposition 4.1 and (4.3), we have
${\mathcal{J}}_{t}^{\lambda}(h_{t})=\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+V_{t}^{2}+\lambda\|\nabla h_{t}\|_{L^{2}(p_{T-t})}^{2}$
Since $h_{t}^{\lambda}$ is the minimizer of ${\mathcal{J}}_{t}^{\lambda}$, the methods of variation imply that for any $v_{t}\in H^{1}(p_{T-t})$,
$\delta{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda},v_{t})=\langle h_{t}^{\lambda}-h_{t}^{*},v_{t}\rangle_{L^{2}(p_{T-t})}+\lambda\langle\nabla h_{t}^{\lambda},\nabla v_{t}\rangle_{L^{2}(p_{T-t})}=0,$
which implies
$\langle h_{t}^{\lambda}-h_{t}^{*},v_{t}\rangle_{L^{2}(p_{T-t})}+\lambda\langle\nabla h_{t}^{\lambda}-\nabla h_{t}^{*},\nabla v_{t}\rangle_{L^{2}(p_{T-t})}$
$=-\lambda\langle\nabla h_{t}^{*},\nabla v_{t}\rangle_{L^{2}(p_{T-t})}$
$=\lambda\langle\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t},v_{t}\rangle_{L^{2}(p_{T-t})},$
where the last equality invokes Lemma B.1. Substituting $v_{t}\coloneq h_{t}^{\lambda}-h_{t}^{*}$ yields
$\|h_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$
$=\lambda\langle\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t},h_{t}^{\lambda}-h_{t}^{*}\rangle_{L^{2}(p_{T-t})}$
(B.8) $\leq\lambda\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}\|h_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})},$
where the last inequality is due to Cauchy-Schwarz inequality. A direct conclusion is
$\|h_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\leq\lambda^{2}\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2}.$
Then plugging this equality into (B.8) yields
$\|\nabla h_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\leq\lambda\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2},$
which completes the proof. ∎
Proposition 4.3.
Let $\lambda>0$, and $h_{t}^{\lambda}$ be the minimizer of ${\mathcal{J}}_{t}^{\lambda}$ defined as (4.3). Then for any $h_{t}\in H^{1}(p_{T-t})$, we have
$\frac{1}{\max\{\lambda,1\}}\big\{{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big\}\leq\|h_{t}-h_{t}^{\lambda}\|_{H^{1}(p_{T-t})}^{2}\leq\frac{1}{\min\{\lambda,1\}}\big\{{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big\}.$
Proof of Proposition 4.3.
Using Proposition 4.1 and (4.3), we have
${\mathcal{J}}_{t}^{\lambda}(h_{t})=\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+V_{t}^{2}+\lambda\|\nabla h_{t}\|_{L^{2}(p_{T-t})}^{2}.$
Since $h_{t}^{\lambda}$ is the minimizer of ${\mathcal{J}}_{t}^{\lambda}$, the methods of variation imply
(B.9) $\delta{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda},v_{t})=\langle h_{t}^{\lambda}-h_{t}^{*},v_{t}\rangle_{L^{2}(p_{T-t})}+\lambda\langle\nabla h_{t}^{\lambda},\nabla v_{t}\rangle_{L^{2}(p_{T-t})}=0,$
for any $v_{t}\in H^{1}(p_{T-t})$. A direct calculation yields
${\mathcal{J}}_{t}^{\lambda}(h_{t})={\mathcal{J}}_{t}^{\lambda}(h_{t}-h_{t}^{\lambda}+h_{t}^{\lambda})$
$=\|h_{t}-h_{t}^{\lambda}+h_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+V_{t}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{\lambda}+\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}$
$={\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})+\|h_{t}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}$
$\quad+2\langle h_{t}-h_{t}^{\lambda},h_{t}^{\lambda}-h_{t}^{*}\rangle_{L^{2}(p_{T-t})}+2\lambda\langle\nabla h_{t}-\nabla h_{t}^{\lambda},\nabla h_{t}^{\lambda}\rangle_{L^{2}(p_{T-t})}$
$={\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})+\|h_{t}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2},$
where the last equality holds from (B.9). This completes the proof. ∎
Appendix C Derivations in Section 5.1
@@
Proposition 5.1.
  1. (i)
    $\underline{B}\leq h_{t}^{*}(\mathbf{x})\leq\bar{B}$;
  2. (ii)
  3. (iii)
Proof of Proposition 5.1.
A direct conclusion of Lemmas C.1, C.4, and C.5. ∎
Appendix D Derivations in Section 5.2
D.1  Oracle inequality of variationally stable Doob’s matching
@@
Suppose Assumptions 1 and 2 hold. Let $t\in(0,T)$ and let ${\mathscr{H}}_{t}$ be a hypothesis class. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequalities hold:
${\mathbb{E}}\Big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]$ $\lesssim\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big\}$
$\quad+\bar{B}^{2}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{\lambda^{2}d\bar{B}^{2}}{\sigma_{T-t}^{8}},$
${\mathbb{E}}\Big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]$ $\lesssim\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{\frac{1}{\lambda}\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}+\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big\}$
$\quad+\frac{\bar{B}^{2}}{\lambda}\Big(\frac{\mathrm{VCdim}({\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{d\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big(\frac{\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{\lambda d\bar{B}^{2}}{\sigma_{T-t}^{8}},$
@@
It follows from Proposition 4.3 and Lemma D.1 that
${\mathbb{E}}\Big[\|\widehat{h}_{t}^{\lambda}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\Big]+\lambda{\mathbb{E}}\Big[\|\nabla\widehat{h}_{t}^{\lambda}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\Big]$
$={\mathbb{E}}\Big[{\mathcal{J}}_{t}^{\lambda}(\widehat{h}_{t}^{\lambda})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\Big]$
$\leq\inf_{h_{t}\in{\mathscr{H}}_{t}}\Big\{\|h_{t}-h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}+\lambda\|\nabla h_{t}-\nabla h_{t}^{\lambda}\|_{L^{2}(p_{T-t})}^{2}\Big\}$